If in a triangle ABC, the exradii r1, r2, r3 are such that r1=2r2=3r3 then prove that the triangle is right angled . Also find a:b:c.

Grade:12

3 Answers

SAGAR SINGH - IIT DELHI
13 years ago
Dear student,
r1=2r2=3r3
Then
r1/6=r2/3=r3/2 =k using ratios
Hence r1=6k,
r2=3k
r3=2k
Hence using the property of solution of trianle the required triangle is equilateral.
Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.
All the best.
Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.
Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

Askiitians Expert
Sagar Singh
B.Tech, IIT Delhi
Nirabhra Agrawal
13 years ago

Dear Sagar ji, 

Will you please solve this question in full as I could not prove the required results with the hints given by you. Few days back I had called and requested you for the same.Thanks.

G P Agrawal

F/o Nirabhra Agrawal

Shashank agrawal
7 years ago
r1=6r by using a property 1/r1+1/r2+1/r3=1/rAs,r1=stanA/2, r=(s-a)tanA/2 ,So,s/(s-a)=6a=5s/6Similarly,b=2s/3,c=s/2,So,triangle is right angled at A